A bullet of mass 25g was moving at the rate of 500ms-1. After passing through a solid block, it continued to move at the rate of 100ms-1. How much work the bullet had to do in passing the block.

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
The work done by the bullet can be given by the change in kinetic energy...
W=K2-K1 =1/2m(V22-V12) =0.5X(.025)*( 5002 – 1002 ) =3000Joules